OpenStudy (liizzyliizz):

find the equation of the tangent line to the curve f(x)= lnx/e^x at x=1

OpenStudy (anonymous):

slope m=f'(1) point (1,f(1)) equation: (y-f(1))=m(x-1)

OpenStudy (turingtest):

^right, so what have you got for f'(x) ?

OpenStudy (liizzyliizz):

wouldn't f'(x) be (1/x - ln(x)) * e^-x ?

OpenStudy (turingtest):

yeah, so what is f'(1) ?

OpenStudy (liizzyliizz):

it would be 1

OpenStudy (liizzyliizz):

?

OpenStudy (turingtest):

\[f'(x)=e^{-x}(\frac1x-\ln x)\]\[f'(1)=e^{-1}(\frac11-\ln 1)=e^{-1}(1-0)=e^{-1}\]

OpenStudy (liizzyliizz):

err I messed up on the e^-1 :/

OpenStudy (turingtest):

so that's your slope, m what's f(1) ?

OpenStudy (liizzyliizz):

0

OpenStudy (anonymous):

right, so now you have: slope m=e^-1 and point (1,0) replace in the equation

OpenStudy (liizzyliizz):

ok so it would be y-0=e^-1(x-1) correct?

OpenStudy (liizzyliizz):

then you would just change the form, which is what im doing now to match the choices in my question

OpenStudy (turingtest):

yeah that's right don't know what for you need it in...

OpenStudy (turingtest):

what form*

OpenStudy (anonymous):

y=e^-1(x-1) is the equation of the tangent line to the curve

OpenStudy (liizzyliizz):

these are the choices- x-ey-1 =0 x+ey-1=0 x-y-1=0 ex+y-1=0 ex-y-1=0

OpenStudy (liizzyliizz):

i feel like when i took this test i did the work somewhat right, but when it came to switching it out i made a careless mistake so i want to know which one was it. :c

OpenStudy (turingtest):

multiply both sides by e I meant :/

OpenStudy (turingtest):

actually something is missing...

OpenStudy (turingtest):

\[\large y=e^{-1}(x-1)\]\[\large ey=x-1\]\[\large x-ey-1=0\]but that's not a choice...

OpenStudy (turingtest):

oh, it is the first one I didn't see that one on the first line

OpenStudy (anonymous):

liizzyliizz the first choice is -x-ey-1 =0? or x-ey-1 =0?

OpenStudy (liizzyliizz):

first choice is x-ey-1 =0

OpenStudy (anonymous):

that is the answer

OpenStudy (liizzyliizz):

well this helped thank you, I know what I did wrong. *sigh*